Blood is accelerated from rest to 29.5 cm/s in a distance of 1.79 cm by the left ventricle of the heart. Determine (a) the acceleration of the blood and (b) the duration of the acceleration.
Answer #1

a]Using third equation of motion,

v^2 = 2gh

0.295^2 = 2*a*1.79e-2

a = 2.431 m/s^2 answer

b] t = v/a = 0.295/2.431 = 0.1213 s answer